OpenStudy (anonymous):

consider the equations f(x)=6x-1 and g(x)=4x^2 select the solution for (f+g)(x)

Directrix (directrix):

(f+g)(x)= f(x) + g(x) = 6x-1 + 4x^2 --> @cookie890 Crank out the sum of these terms. Write then is descending powers of x. So, (f+g)(x)= 4x^2 + ? + ?